# Aligning Information Capacity Between Vision and Language via Dense-to-Sparse Feature Distillation for Image-Text Matching

### v1

@@ 1 Introduction
However, most traditional Visual Semantic Embedding (VSE) models overlook the inherent differences between image and text modalities [3, 35, 18]. Specifically, visual signals represent natural, objective records of real-world phenomena, whereas textual signals are inherently semantic and subjective, shaped by human interpretation. Consequently, the same image may be described in various ways depending on individual perspectives. Existing studies refer to this issue as the “ambiguity issue” [36, 4], “sparse annotations” [15] and “multi-view descriptions” [20, 33]. But, at its core, the problem stems from the inherent differences between image and text modalities, particularly the disparity in “information density” [26].
@@ 5 Conclutions
Supplementary Material
6 Additional Implementation Details
Details of Dense Text Generation. For dense text generation, we use LLaVa [23] to generate dense textual descriptions. The hyperparameters for LLaVa are set as follows: Top-P $=0.9$, temperature$=0.2$, and max-new-tokens$=500$. We generate a dense text description for each image in the MS-COCO and Flickr30K datasets, with each dense text sentence corresponding to $\frac{1}{5}$ of the sparse text descriptions. As shown in Table 6, we compare the generated image-dense text dataset with the original dataset. Note that, we only generate the dense texts for training dataset.
Dataset Images Texts Tokens per Text
Flickr30K original 29000 145000 13.41
Flickr30K generated 29000 29000 173.24
MS-COCO original 113287 566435 11.30
MS-COCO generated 113287 113287 174.34
Table 6: Comparison between the generated data with the original datasets.
Model settings. The dimensionality $d$ of the joint embedding space is 512 for all experiments. For the image encoder, we utilize different backbones to extract image features and employ a feature aggregator to obtain a holistic embedding. Specifically, we adopt GPO [3] as the feature aggregator. Since different backbones produce features with varying dimensionalities, we introduce an MLP before GPO to standardize the feature dimensions. Similarly, for the text encoder, we use different backbones along with GPO for feature aggregation. For the text decoder, we adopt a Transformer-based architecture with a depth of 4 and 4 attention heads per layer. Additionally, we introduce 100 learnable tokens to predict the context of sparse text embeddings. Notably, the image encoder and text encoder maintain the same architecture across both the pre-training and fine-tuning stages.
Algorithm 1: PyTorch Pseudocode of Finetune Stage
# n: batch_size # d: dimensionality of the embeddings # m: number of the words in sentences # i_e: image embedding # st_e: sparse text embedding # st_w: sparse text word embedding # dt_e: dense text embedding # compute embeddings i_e = image_encoder(I) #[n, d] st_e, st_w = sparse_text_encoder(T) # [n, d] [n,m,d] st_e = decoder(st_w) + st_e with torch.no_grad(): dt_e, _ = dense_text_encoder(T) # [n,d] # normalize i_e_norm = l2_normalize(i_e, dim=1) # [n,d] st_e_norm = l2_normalize(st_e, dim=1) # [n,d] dt_e_norm = l2_normalize(dt_e, dim=1) # [n,d] # contrastive learning sims = st_e_norm.mm(i_e_norm.t()) loss_align = triplet(sims) # d2s distill loss_distill = distillation_loss(dt_e_norm, st_e_norm) loss = loss_align + loss_distill
Training details. For dense text pretraining, the VSE model is optimized using AdamW with a weight decay factor of 1e-4 The batch size is set to 96 for a Swin-based backbone and 128 for other architectures. The margin $\alpha$ of the triplet loss is 0.2. The initial learning rate is set to 0.0005 for the first few epochs and is progressively reduced by a factor of 10 at the 9th, 15th, 20th, and 25th epochs. Additionally, a warm-up strategy is applied at the beginning of training: during the first epoch, the triplet loss considers all negative examples in the batch instead of selecting only the hardest negative example.For sparse text fine-tuning, we initialize the image and text encoders with weights from the pretraining phase. The objective of fine-tuning is to reconstruct the dense text embedding from the sparse text embedding. The training strategy remains consistent with the pretraining phase.
The pseudo-code for the fine-tuning phase is presented in Algorithm 1.

### v2

@@ 1 Introduction
However, most traditional Visual Semantic Embedding (VSE) models overlook the inherent differences between image and text modalities [3, 35, 18]. Specifically, visual signals represent natural, objective records of real-world phenomena, whereas textual signals are inherently semantic and subjective, shaped by human interpretation. Consequently, the same image may be described in various ways depending on individual perspectives. Existing studies refer to this issue as the “ambiguity issue" [36, 4], “sparse annotations" [15] and “multi-view descriptions" [20, 33]. But, at its core, the problem stems from the inherent differences between image and text modalities, particularly the disparity in “information density" [26].